OpenStudy (mikezack123):

absolute value! check my work please! |2t-3| = 3t-2 i got case 1 t=-1 extraneous case 2 t= 1 extraneous so my answer is no solution

OpenStudy (mikezack123):

@myininaya

OpenStudy (mikezack123):

@zepdrix @dan815 @jim_thompson5910

OpenStudy (aum):

t = -1 and t = 1 are correct. If we try t = -1 we get LHS = |-2-3| = 5 and RHS = -3-2 = -5. (extraneous) If we try t = 1 we get LHS = |2-3| = 1 and RHS = 3-2 = 1. t = 1 is a solution.

OpenStudy (mikezack123):

okay so only case 2 works? and case 1 is extraneous?

OpenStudy (aum):

Yes.

OpenStudy (mikezack123):

okay thank you @aum

OpenStudy (aum):

You are welcome.

jimthompson5910 (jim_thompson5910):

visually you can compare the graphs of y = |2x-3| and y = 3x-2 wherever they intersect is where your solution will be turns out they only intersect at one point: (1,1) x = 1 ---> t = 1 is your only solution see attached for the graph

myininaya (myininaya):

@MikeZack123 When you check those solutions, make sure both sides are the same (if they are both the same it is a solution) If both sides are not the same when pluggin (testing your number) in your number, then it is extraneous (or not a solution)

OpenStudy (mikezack123):

yes thank you @jim_thompson5910 and @myininaya :) much appreciated